OpenStudy (anonymous):

A mass m1 = 5.5 kg rests on a frictionless table and connected by a massless string to another mass m2 = 4.6 kg. A force of magnitude F = 37.0 N pulls m1 to the left a distance d = 0.7 m. What is the final speed of the two blocks? How much work is done by the tension (in-between the blocks) on block m2? What is the tension in the string? What is the NET work done on m1?

OpenStudy (anonymous):

Work = F*d = delta KE since they presumably start at V=0 F*d = 1/2 * 5.5*V^2 + 1/2*4.6*V^2 calculate F*d and use it to solve for V

OpenStudy (anonymous):

for the second and third parts F*d - 1/2*5.5*V^2 = work done on the second block You can use that to find the tension, since the distance is the same for both blocks, so: (Force of tension in the string between m1 and m2)*d = F*d - 1/2*5.5*V^2

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

yes thank you!